the measure of an angles supplement is 44 less than the measure of the angle. Find the measure of the angle and it’s supplement.

Answer by ikleyn(53937)

.
the measure of an angles supplement is 44 less than the measure of the angle.
Find the measure of the angle and it’s supplement.
~~~~~~~~~~~~~~~

Let x be the measure of the angle; then the supplementary angle is (180-x) degrees,

and the problem gives us this equation to find x


    x - (180-x) = 44  degrees.


Simplify and find x


    2x = 44 + 180 = 224

     x            = 224/2 = 112.


ANSWER.  The measure of the angle is 112 degrees;  the measure of its supplement is 180 - 112 = 68.


CHECK.   The difference of the measures is  112 - 68 = 44  degrees.   ! Correct !

Solved.